February 28th 2025
From Hall A Wiki
- Updates
- CDet
- internal cables being installed, after cables to repeater , aiming to have everything cabled by Monday
- need install scintillator trigger
- will takes cosmics next week
- GEM
- CODA was working, few APVs have issues, being checked
- Paolo : Marco will bring MPDs March 8th
- Inventory of MPD by Ching
- 3 MPDs given to Ben : 2 MPDs out of 3 working - one MPD with issue with Digital output
- Rafael : Anu gave sample GEM layers X an Y region of interest and outputting APV and MPD corresponding
- CDet